# The Interplay of Heterogeneity and Product Detachment in Templated Polymer Copying

### v1

@@ II.4 Recovering Further Properties of the Fine-Grained System from the Coarse-Grained System
Here, $\mathbb{E}[V_{\mathbf{x},l}(m_{l},m_{l+1})|m_{l-1},m_{l}]$ is the expected number of visits to tip state $(\mathbf{x},m_{l}m_{l+1},l+1)$ from each visit of $(\mathbf{x},m_{l-1}m_{l},l)$. $\mathbb{E}[V_{\mathbf{x},l}(m_{l},m_{l+1})|m_{l-1},m_{l}]$ can be obtained from the Markov chain constructed in Figure 3. This Markov chain is similar to that in Figure 2 with two important changes. First, instead of considering absorption probabilities from $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$, we further extend the chain to $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1}m_{l+2})$ before allowing absorption (these absorption probabilities are again the $Q$ probabilities from Gaspard’s iteration in Section II.3), allowing visits to $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$ from $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1}m_{l+2})$ to be counted. Second, monomer removal from $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$ results in an artificial absorbing state instead of returning to $(\mathbf{x},\&m_{l-1}m_{l})$. This approach prevents the double counting of visits to $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$ from distinct visits of $(\mathbf{x},\&m_{l-1}m_{l})$. Note that we have not imposed any additional assumptions unique to our system for this iteration to work, and it is applicable to any polymerisation system whose coarse-grained form (applying Qureshi’s procedure [43]) obeys the assumptions in [32]. The derivation of a general form for $\mathbb{E}[V_{\mathbf{x},l+1}(m_{l}m_{l+1})|m_{l-1}m_{l}]$ is provided in Appendix D.
Figure 3: Markov chain from which expected visitation counts per visit of the previous tip state $\mathbb{E}[V_{\mathbf{x},l}(m_{l},m_{l+1})|m_{l-1},m_{l}]$ can be calculated. Solid arrows represent local transition probabilities that can be calculated from coarse-grained rates $\Phi^{+}$ or $\Phi^{-}$, while dotted arrows represent the $Q$ variables obtained via Gaspard’s iteration (Figure 2). To avoid over-counting, the tip states whose visitation counts are of interest are not permitted to move back into the initial state, rather they point to a virtual absorbing state at a rate equal to the rate of monomer removal. Starting from the initial state, we can now count, through standard methods, the average number of times either tip state of interest is visited before either the initial tip state is revisited or polymerization terminates.
@@ II.5 Thermodynamically Consistent Parameterization for Heterogeneous Separating Copiers
Now that the fine-grained system has been parameterized, we can proceed to define transition propensities [43]. These propensities are derived in full in Appendix B, and the results are given in equations 8 and 9. Due to the forms taken by our fine-grained steps, we can impose the stronger condition on locality than required by Gaspard’s method: that the propensities are also dependent only on local template monomers, as well as copy monomers, $\Phi^{+}(m_{l+1},\mathbf{x},\&m_{l-1}m_{l})=\Phi^{+}(n_{l}n_{l+1},m_{l}m_{l+1})$ and $\Phi^{-}(n_{l-1}n_{l},\&m_{l-1}m_{l})=\Phi^{-}(n_{l-1}n_{l},m_{l-1}m_{l})$. We find
@@ III.1 Consequences of Separation for the Velocity and Energy Landscapes of Heterogeneous Copying
We wish to compare the velocity profile of heterogeneous copying (with separation) and templated self-assembly (without separation, as studied by Gaspard [32]; abbreviated as TSA). In figure 4(a), we illustrate a fine-grained model of templated self-assembly which we will compare against our model of templated copying. The fine-grained dynamics is similar to that in Section II.5, with the exception that the tail binding step has been removed. We assume here that the binding free energy of any given monomer pair is independent of any other monomer pair, giving rise to the parameterization in Table 3. We consider here the slow binding limit, leading to the following coarse-grained propensities:
@@
This choice again corresponds to discrimination on backward propensities. We denote the sequence-independent non-equilibrium drive provided by monomer binding and backbone formation by $\Delta G_{\textnormal{pol}}=\Delta G_{BB}+\ln{[M]}$ (similar to copying with separation, but with the $[M]_{\textnormal{eff}}$ term absent). However, observe that in the absence of separation, the sequence-dependent free-energy change $\Delta G_{PT}$ incurred on the addition of each monomer persists as the copy is produced.
@@
We begin by considering how the velocity profiles for heterogeneous templated copying change with increasing nonequilibrium drive $\Delta G_{\textnormal{pol}}$. As with TSA, we consider backward propensity discrimination (that is, the slow binding limit). The binding free energies of the incorrect monomer (we remind readers that a correct pair is defined by $m_{l}=n_{l}$) pairs are kept constant, such that $\Delta G_{TT}(1,2)=\Delta G_{TT}(2,1)=0$. The free energy of a correct monomer $1$ pair is held at $\Delta G_{TT}(1,1)=2$, while $\Delta G_{TT}(2,2)$ is varied from $2$ to $10$. Gaspard’s iterated function system [32] was applied to obtain the absorption probabilities $Q$ for the copying of a polymer sequence of length $L=10000$. Then, the methods in Section II.4 and reference [43] were used to find the average completion times $t_{c}$, and the average copying velocity was calculated as $v=\frac{L}{t_{c}}$. As in reference [33], the value of $\Delta G_{\textnormal{pol}}$ at equilibrium is given by $\Delta G_{\textnormal{pol,eq}}=-\ln{2}$, and we report copy velocities as a function of the difference in $\Delta G_{\textnormal{pol}}$ from this equilibrium value up to $\Delta G_{\textnormal{pol}}=2$ in figure 5(b).
@@
Figure 4: Fine-grained steps for a model of copying where the copy does not separate from the template. (a) New monomers bind and polymerize in separate steps, but there is no tail unbinding of the previously added monomer. (b) Labelling of fine-grained states associated with a given coarse-grained state. Indices $f$ are denoted by bold bracketed numbers. Transitions to other states of index $f=0$ (denoted here by the dotted arrows) change the coarse-grained state $(\mathbf{x},\mathbf{y})$.
@@
Figure 5: Comparison of velocity profiles and free energy landscapes for copying with separation and TSA without separation. Graphs of equilibrium velocity as a function of $\Delta G_{\textnormal{pol}}-\Delta G_{\textnormal{pol,eq}}$ are shown for TSA in (a) and copying with separation in (b). Equilibrium $G_{\mathbf{x},l}-\langle G\rangle$ free-energy landscapes are shown for TSA in (c) and copying with separation (d). A zoom inset is provided in (d). The long tails of zero velocity for copying without separation, caused by a rougher free-energy profile, do not occur for copying with simultaneous separation, where the velocity profile is linear close to $\Delta G_{\textnormal{pol}}-\Delta G_{\textnormal{pol,eq}}=0$.
@@
We now consider velocity profiles for TSA without separation. We must first calculate the equilibrium point $\Delta G_{\textnormal{pol,eq}}$, which follows from the partition function of the product. As the monomer binding free energies are uncorrelated in our TSA system, the (negative of) free energy $G_{\mathbf{x},l}$ of the ensemble of length of length $l$ product polymers can be calculated as follows:
@@
We plot the normalized copy velocity (obtained by dividing velocity by its maximum value over the range of $\Delta G_{\textnormal{pol}}$ considered; in practice this maximum value always occurs at our maximum considered drive $\Delta G_{\textnormal{pol}}=0$) $\overline{v}=v/(\max_{\Delta G_{\textnormal{pol}}}v)$ as a function of $\Delta G_{\textnormal{pol}}-\Delta G_{\textnormal{pol,eq}}$ for TSA in Figure 5(a). Comparing Figures 5(a) and 5(b), we observe a few common features. First, velocity tends to zero as the equilibrium $\Delta G_{\textnormal{pol}}$ is reached. At equilibrium there should be no net flux, and hence this behaviour is expected. Copy velocity increases with increasing $\Delta G_{\textnormal{pol}}$, as the rate of monomer removal decreases.
However, the models approach zero velocity in different ways. As noted by Gaspard [32], the heterogeneous model of templated self-assembly displays a long tail of near-zero velocity as it approaches equilibrium. Such behaviour is not observed in our model of heterogeneous copying (Figure 5(b)). This difference can be explained by considering sample free-energy landscapes in the heterogeneous TSA model and the heterogeneous copying model. We plot free-energy deviation $G_{\mathbf{x},l}-\langle G\rangle$ for a fixed sample template at various copy lengths $l$ for both heterogeneous TSA (Figure 5(c)) and heterogeneous copying (Figure 5(d)).
For heterogeneous copying, the free-energy deviations switch stochastically between two levels, determined by the template monomer at position $l$, and $G_{\mathbf{x},l}-\langle G\rangle$ is therefore constrained to be close to zero. By contrast, the free-energy deviations for heterogeneous TSA are large. On scales of length $l_{d}$ we expect to see barriers of height $\sim\sqrt{l_{d}}$ [47] that trap the system, resulting in slow sub-diffusive motion close to equilibrium driving [47, 32]. Hence we observe a region of zero velocity in Figure 5(a). These tall barriers do not occur for separating copiers as the length of product interacting with the template at any given time is bounded. In our system, only two monomers may be bound at a given time, but this intuition generalizes to other parameter sets and even completely different models that include separation as the copy grows (even realistic models of transcription and translation). If only a finite number of copy monomers bind to the template simultaneously, the roughness of the free-energy landscape associated with sequence heterogeniety is inherently limited.
@@ III.2.3 Discussion
Heterogeneity on the correct monomers tends to make copying both slower (up to modifications due to waiting time) and more error-prone, while heterogeneity on incorrect monomers tends to make copying faster and more accurate (in some cases, only the average relative error is made better). We now attempt to explain these results. We can divide the overall coarse-grained state-space into two: one in which the correct monomer is bound at the tip, and one in which the incorrect monomer is bound at the tip. Due to detachment of the copy behind the tip, this division is enough to specify the chemical free energy at each step (the same does not apply to models of TSA). Copying can then be thought of as a special example of 1D diffusion with a choice of free-energy landscape at each step (Figure 9). The addition of a correct monomer corresponds to moving through a more favourable landscape (the ‘correct landscape’), while the addition of an incorrect monomer involves moving through a less favourable landscape (the ‘incorrect landscape’). To understand the effects of heterogeneity, we only need to consider transitions with free-energy changes modified by heterogeneity (purple lines in Figure 9), as all other transitions would be found in equivalent forms in corresponding homogeneous copying systems. Observe that the introduction of heterogeneity in the interaction of correct monomers results in an increased roughness of the correct landscape, while the introduction of heterogeneity in the interaction of incorrect monomers results in an increased roughness of the incorrect landscape. This relative roughness is particularly evident as the overall variability in $\Delta G_{\mathbf{x},l}$ is constrained within finite bounds (Figure 5(d)).
@@
We perform a sweep over parameters to identify regimes where heterogeniety results in the greatest improvement in error rate (results in Appendix F). Interestingly, our parameter sweep revealed parameter sets where a heterogeneous system has a lower mean error than either of its constituent monomers used in isolation. Consider for instance the plot in Figure 10. There is a clear minimum in $\langle\epsilon\rangle$ at $p_{t}=0.4$, implying that some combination of the two considered monomers performs better than either one on its own. Thus, it is generally untrue that the error performance of a heterogeneous copying system is bounded by the error performance of its constituent monomers.
@@ IV Conclusion
Using minimal models with fine-grained steps [35], we have investigated how heterogeneity interacts with separation to affect copying error, velocity and thermodynamiv efficiency in templated copying. Our first contribution is an approach for extending Qureshi’s coarse-graining method [43] to generalize to heterogeneous systems. We have thus far used this method to find polymer completion times for our coarse-grained model. A natural extension would be to use the method to investigate a heterogeneous version of kinetic proofreading [36, 43], where it can be used to find the average expected free-energy consumption per polymer copy.
Using this method, we were able to characterize velocity profiles of heterogeneous copying systems that simultaneously separate. In contrast to templated self-assembly without separation [32], we do not observe a long tail of zero velocity as equilibrium is approached. This absence makes sense in light of the template-dependent free-energy profiles of partial copies, which have significantly higher barriers (scaling as $\sqrt{l_{D}}$ for a length scale $l_{D}$) in the case of TSA. Put another way, each monomer type in heterogeneous TSA has a different pseudo-equilibrium point, and copying with a drive lower than a monomer’s pseudo-equilibrium point makes traversing a long stretch of said monomer more difficult. This effect is totally absent in the case of copying with simultaneous separation, since the sequence-specific interactions with the template are transient. We expect that this observation would generalise to more realistic models of transcription or translation, as long as the copy continuously separates from the template and the number of copy monomers interacting with the template at a given time is effectively bounded.

### v2

@@ II.4 Recovering Further Properties of the Fine-Grained System from the Coarse-Grained System
Here, $\mathbb{E}[V_{\mathbf{x},l}(m_{l},m_{l+1})|m_{l-1},m_{l}]$ is the expected number of visits to tip state $(\mathbf{x},m_{l}m_{l+1},l+1)$ from each visit of $(\mathbf{x},m_{l-1}m_{l},l)$. $\mathbb{E}[V_{\mathbf{x},l}(m_{l},m_{l+1})|m_{l-1},m_{l}]$ can be obtained from the Markov chain constructed in Figure 3. This Markov chain is similar to that in Figure 2 with two important changes. First, instead of considering absorption probabilities from $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$, we further extend the chain to $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1}m_{l+2})$ before allowing absorption (these absorption probabilities are again the $Q$ probabilities from Gaspard’s iteration in Section II.3), allowing visits to $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$ from $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1}m_{l+2})$ to be counted. Second, monomer removal from $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$ results in an artificial absorbing state instead of returning to $(\mathbf{x},\&m_{l-1}m_{l})$. This approach prevents the double counting of visits to $(\mathbf{x},\&m_{l-1}m_{l}m_{l+1})$ from distinct visits of $(\mathbf{x},\&m_{l-1}m_{l})$. Note that we have not imposed any additional assumptions unique to our system for this iteration to work, and it is applicable to any polymerisation system whose coarse-grained form (applying Qureshi et al.’s procedure [43]) obeys the assumptions in [32]. The derivation of a general form for $\mathbb{E}[V_{\mathbf{x},l+1}(m_{l}m_{l+1})|m_{l-1}m_{l}]$ is provided in Appendix D.
Figure 3: Markov chain from which expected visitation counts per visit of the previous tip state $\mathbb{E}[V_{\mathbf{x},l}(m_{l},m_{l+1})|m_{l-1},m_{l}]$ can be calculated. Solid arrows represent local transition probabilities that can be calculated from coarse-grained rates $\Phi^{+}$ or $\Phi^{-}$, while dotted arrows represent the $Q$ variables obtained via Gaspard’s iteration [32] (Figure 2). To avoid over-counting, the tip states whose visitation counts are of interest are not permitted to move back into the initial state, rather they point to a virtual absorbing state at a rate equal to the rate of monomer removal. Complete polymers are again treated as absorbing states. Starting from the initial state, we can now count, through standard methods, the average number of times either tip state of interest (denoted by “$\textnormal{Tip}(m_{l},m_{l}+1)_{l}$”) is visited before either the initial tip state is revisited or polymerization terminates.
@@ II.5 Thermodynamically Consistent Parameterization for Heterogeneous Separating Copiers
Now that the fine-grained system has been parameterized, we can proceed to define transition propensities [43]. These propensities are derived in full in Appendix B, and the results are given in equations 8 and 9. Due to the forms taken by our fine-grained steps, we can impose a more restrictive condition on locality than required by Gaspard’s method: that the propensities are also dependent only on local template monomers, as well as copy monomers, $\Phi^{+}(m_{l+1},\mathbf{x},\&m_{l-1}m_{l})=\Phi^{+}(n_{l}n_{l+1},m_{l}m_{l+1})$ and $\Phi^{-}(n_{l-1}n_{l},\&m_{l-1}m_{l})=\Phi^{-}(n_{l-1}n_{l},m_{l-1}m_{l})$. We find
@@ III.1 Consequences of Separation for the Velocity and Energy Landscapes of Heterogeneous Copying
We wish to compare the velocity profile of separating templated copolymerization (STC) and non-separating templated copolymerization (NTC, note the phrase ‘templated self-assembly’ has been used in some literature [33]) as studied by Gaspard [32]. In figure 4(a), we illustrate a fine-grained model of NTC which we will compare against our model of STC. The fine-grained dynamics is similar to that in Section II.5, with the exception that the tail binding step has been removed. We assume here that the binding free energy of any given monomer pair is independent of any other monomer pair, giving rise to the parameterization in Table 3. We consider here the slow binding limit, leading to the following coarse-grained propensities:
@@
This choice again corresponds to discrimination on backward propensities. We denote the sequence-independent non-equilibrium drive provided by monomer binding and backbone formation by $\Delta G_{\textnormal{pol}}=\Delta G_{BB}+\ln{[M]}$ (similar to copying with separation, but with the $[M]_{\textnormal{eff}}$ term absent). However, observe that in the absence of separation, the sequence-dependent free-energy change $\Delta G_{PT}$ incurred on the addition of each monomer persists as the copy is produced (note $\Delta G_{TT}$ was used in [33] to refer to a ‘temporary thermodynamic’ discrimination factor, that is no longer present after separation. To maintain consistency with the intended meaning in [33], we have introduced a ‘permanent thermodynamic’ discrimination factor $\Delta G_{PT}$ for NTC).
@@
We begin by considering how the velocity profiles for STC change with increasing nonequilibrium drive $\Delta G_{\textnormal{pol}}$. As with NTC, we consider backward propensity discrimination (that is, the slow binding limit). The binding free energies of the incorrect monomer (we remind readers that a correct pair is defined by $m_{l}=n_{l}$) pairs are kept constant, such that $\Delta G_{TT}(1,2)=\Delta G_{TT}(2,1)=0$. The free energy of a correct monomer $1$ pair is held at $\Delta G_{TT}(1,1)=2$, while $\Delta G_{TT}(2,2)$ is varied from $2$ to $10$. Gaspard’s iterated function system [32] was applied to obtain the absorption probabilities $Q$ for the copying of a polymer sequence of length $L=10000$. Then, the methods in Section II.4 and reference [43] were used to find $\tau_{\mathbf{x},l}$, defined as the average amount of total time spent at position $l$ for a template sequence $\mathbf{x}$ before a complete polymer forms and detaches. $\tau_{\mathbf{x},l}$ is calculated by multiplying visits to each state by the waiting time for each visit to said state and then summing over all states of copy length $l$. The average completion times $t_{c}=\Sigma_{l}\tau_{\mathbf{x},l}$ can then be calculated, and the average copying velocity was calculated as $v=\frac{L}{t_{c}}$. As in reference [33], the value of $\Delta G_{\textnormal{pol}}$ at equilibrium is given by $\Delta G_{\textnormal{pol,eq}}=-\ln{2}$, and we report copy velocities as a function of the difference in $\Delta G_{\textnormal{pol}}$ from this equilibrium value up to $\Delta G_{\textnormal{pol}}=2$ in figure 5(b).
@@
Figure 4: Fine-grained steps for a model of copying where the copy does not separate from the template. Only monomers in the vicinity of the copy tip are depicted. (a) New monomers bind and polymerize in separate steps, but there is no tail unbinding of the previously added monomer. (b) Labelling of fine-grained states associated with a given coarse-grained state. Indices $f$ are denoted by bold bracketed numbers. Transitions to other states of index $f=0$ (denoted here by the dotted arrows) change the coarse-grained state $(\mathbf{x},\mathbf{y})$.
@@
Figure 5: Comparison of velocity profiles and free energy landscapes for STC and NTC. Graphs of normalized average velocity as a function of $\Delta G_{\textnormal{pol}}-\Delta G_{\textnormal{pol,eq}}$ are shown for NTC in (a) and STC in (b). Equilibrium $G_{\mathbf{x},l}-\langle G\rangle$ free-energy landscapes (at $\Delta G_{pol}=\Delta G_{pol,eq}$) are shown for NTC in (c) and STC (d). A zoom inset is provided in (d). The long tails of zero velocity for copying without separation, caused by a rougher free-energy profile, do not occur for copying with simultaneous separation, where the velocity profile is linear close to $\Delta G_{\textnormal{pol}}-\Delta G_{\textnormal{pol,eq}}=0$.
@@
We now consider velocity profiles for NTC. We must first calculate the equilibrium point $\Delta G_{\textnormal{pol,eq}}$, which follows from the partition function of the product. As the monomer binding free energies are uncorrelated in our NTC system, the (negative of) free energy $G_{\mathbf{x},l}$ of the ensemble of length of length $l$ product polymers can be calculated as follows:
@@
We plot the normalized copy velocity (obtained by dividing velocity by its maximum value over the range of $\Delta G_{\textnormal{pol}}$ considered; in practice this maximum value always occurs at our maximum considered drive $\Delta G_{\textnormal{pol}}=0$) $\overline{v}=v/(\max_{\Delta G_{\textnormal{pol}}}v)$ as a function of $\Delta G_{\textnormal{pol}}-\Delta G_{\textnormal{pol,eq}}$ for NTC in Figure 5(a). Comparing Figures 5(a) and 5(b), we observe a few common features. First, velocity tends to zero as the equilibrium $\Delta G_{\textnormal{pol}}$ is reached. At equilibrium there should be no net flux, and hence this behaviour is expected. Copy velocity increases with increasing $\Delta G_{\textnormal{pol}}$, as the rate of monomer removal decreases.
However, the models approach zero velocity in different ways. As noted by Gaspard [32], the heterogeneous model of NTC displays a long tail of near-zero velocity as it approaches equilibrium. Such behaviour is not observed in our model of heterogeneous copying (Figure 5(b)). This difference can be explained by considering sample free-energy landscapes in the heterogeneous NTC model and the STC model. We plot free-energy deviation $G_{\mathbf{x},l}-\langle G\rangle$ for a fixed sample template at various copy lengths $l$ for both heterogeneous NTC (Figure 5(c)) and heterogeneous STC (Figure 5(d)).
For heterogeneous STC, the free-energy deviations switch stochastically between two levels, determined by the template monomer at position $l$, and $G_{\mathbf{x},l}-\langle G\rangle$ is therefore constrained to be close to zero. By contrast, the free-energy deviations for heterogeneous NTC are large. On scales of length $l_{d}$ we expect to see barriers of height $\sim\sqrt{l_{d}}$ [50] that trap the system, resulting in slow sub-diffusive motion close to equilibrium driving [50, 32]. Hence we observe a region of zero velocity in Figure 5(a). These tall barriers do not occur for separating copiers as the length of product interacting with the template at any given time is bounded. In our system, only two monomers may be bound at a given time, but this intuition generalizes to other parameter sets and even completely different models that include separation as the copy grows (even realistic models of transcription and translation). If only a finite number of copy monomers bind to the template simultaneously, the roughness of the free-energy landscape associated with sequence heterogeneity is inherently limited.
@@ III.2.3 Discussion
Heterogeneity on the correct monomers tends to make copying both slower (up to modifications due to waiting time) and more error-prone, while heterogeneity on incorrect monomers tends to make copying faster and more accurate (in some cases, only the average relative error is made better). We now attempt to explain these results. We can divide the overall coarse-grained state-space into two: one in which the correct monomer is bound at the tip, and one in which the incorrect monomer is bound at the tip. Due to detachment of the copy behind the tip, this division is enough to specify the chemical free energy at each step (the same does not apply to models of NTC). Copying can then be thought of as a special example of 1D diffusion with a choice of free-energy landscape at each step (Figure 9). The addition of a correct monomer corresponds to moving through a more favourable landscape (the ‘correct landscape’), while the addition of an incorrect monomer involves moving through a less favourable landscape (the ‘incorrect landscape’). To understand the effects of heterogeneity, we only need to consider transitions with free-energy changes modified by heterogeneity (purple lines in Figure 9), as all other transitions would be found in equivalent forms in corresponding homogeneous copying systems. Observe that the introduction of heterogeneity in the interaction of correct monomers results in an increased roughness of the correct landscape, while the introduction of heterogeneity in the interaction of incorrect monomers results in an increased roughness of the incorrect landscape. This relative roughness is particularly evident as the overall variability in $\Delta G_{\mathbf{x},l}$ is constrained within finite bounds (Figure 5(d)).
@@
We perform a sweep over parameters to identify regimes where heterogeneity results in the greatest improvement in error rate (results in Appendix F). Interestingly, our parameter sweep revealed parameter sets where a heterogeneous system has a lower mean error than either of its constituent monomers used in isolation. Consider for instance the plot in Figure 10. There is a clear minimum in $\langle\epsilon\rangle$ at $p_{t}=0.4$, implying that some combination of the two considered monomers performs better than either one on its own. Thus, it is generally untrue that the error performance of a heterogeneous copying system is bounded by the error performance of its constituent monomers.
@@ IV Conclusion
Using minimal models with fine-grained steps [35], we have investigated how heterogeneity interacts with separation to affect copying error, velocity and thermodynamic efficiency in STC. We have thus far investigated heterogeneity in monomer binding energies and kinetics; the effects of heterogeneous monomer concentrations, which would naturally manifest as heterogeneity in binding rates, will be an interesting topic for future research. Our first contribution is an approach for extending Qureshi et al’s coarse-graining method [43] to generalize to heterogeneous systems. We have thus far used this method to find polymer completion times for our coarse-grained model. A natural extension would be to use the method to investigate a heterogeneous version of kinetic proofreading [36, 43], where it can be used to find the average expected free-energy consumption per polymer copy.
Using this method, we were able to characterize velocity profiles of a heterogeneous separating templated copolymerization system. In contrast to non-separating templated copolymerization systems [32], we do not observe a long tail of zero velocity as equilibrium is approached. This absence makes sense in light of the template-dependent free-energy profiles of partial copies, which have significantly higher barriers (scaling as $\sqrt{l_{D}}$ for a length scale $l_{D}$) in the case of NTC. Put another way, each monomer type in heterogeneous NTC has a different pseudo-equilibrium point, and copying with a drive lower than a monomer’s pseudo-equilibrium point makes traversing a long stretch of said monomer more difficult. This effect is totally absent in the case of STC, since the sequence-specific interactions with the template are transient. We expect that this observation would generalise to more realistic models of transcription or translation, as long as the copy continuously separates from the template and the number of copy monomers interacting with the template at a given time is effectively bounded.